OpenStudy (anonymous):

Graph the following linear inequality. 6x+5y<5

OpenStudy (anonymous):

Convert to y<mx + b Then graph it like normal. Test the point on either side of the graph you just created, if it doesn't make sense then the region which you pick the point is not true, so it has to be the other one. That's it! That's all you need to graph it. Try this and see if you get the right answer :)

OpenStudy (anonymous):

5y<6x+5?

OpenStudy (anonymous):

5y< -6x + 5 then divide both sides by 5, you get y<-6x/5 + 5

OpenStudy (anonymous):

Graph that :)

OpenStudy (anonymous):

\[y<\frac{ -6x }{ 5 }+5?\]

OpenStudy (anonymous):

Or does the x belong on the outside of the fraction?

OpenStudy (anonymous):

outside, or inside, it's the same thing :P

OpenStudy (anonymous):

oh, and one more thing. Since this inequality is a "less than", remember to draw your graph with dash lines

OpenStudy (anonymous):

Got it! So I just take random #s and sub them for x? like -1,0,1?

OpenStudy (anonymous):

both x and y, just plug it in, see if it makes any sense

OpenStudy (anonymous):

Ok, ill try

OpenStudy (anonymous):

I'm lost, can you find one point cuz im pretty sure im doing it wrong

OpenStudy (anonymous):

tell me what you did

OpenStudy (anonymous):

-1<-6/5(-1)+5

OpenStudy (anonymous):

so... is that true?

OpenStudy (anonymous):

the right hand side is 6.2, so obviously 6.2 is greater than -1 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so because that region that you tested give you the right solution, should should color that side in

OpenStudy (anonymous):

and you're done!

OpenStudy (anonymous):

how did u get 6.2?

OpenStudy (anonymous):

-6/5 * (-1) +5 = 6/5 + 5 = 1.2 + 5 = 6.2

OpenStudy (anonymous):

Ohhhhhh, ok so I plot -1,-1?

OpenStudy (anonymous):

no, you just use a point on the region of the graph to see whether if your given equation decribes that region or not

OpenStudy (anonymous):

So what do I plot?

OpenStudy (anonymous):